\begin{thebibliography}{74}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Adiwardana et~al.(2020)Adiwardana, Luong, So, Hall, Fiedel,
  Thoppilan, Yang, Kulshreshtha, Nemade, Lu et~al.}]{adiwardana2020towards}
Daniel Adiwardana, Minh-Thang Luong, David~R So, Jamie Hall, Noah Fiedel, Romal
  Thoppilan, Zi~Yang, Apoorv Kulshreshtha, Gaurav Nemade, Yifeng Lu, et~al.
  2020.
\newblock Towards a human-like open-domain chatbot.
\newblock \emph{arXiv preprint arXiv:2001.09977}.

\bibitem[{Alamri et~al.(2019)Alamri, Cartillier, Das, Wang, Cherian, Essa,
  Batra, Marks, Hori, Anderson et~al.}]{alamri2019audio}
Huda Alamri, Vincent Cartillier, Abhishek Das, Jue Wang, Anoop Cherian, Irfan
  Essa, Dhruv Batra, Tim~K Marks, Chiori Hori, Peter Anderson, et~al. 2019.
\newblock Audio visual scene-aware dialog.
\newblock In \emph{Proceedings of the IEEE Conference on Computer Vision and
  Pattern Recognition}, pages 7558--7567.

\bibitem[{Asghar et~al.(2018)Asghar, Poupart, Hoey, Jiang, and
  Mou}]{asghar2018affective}
Nabiha Asghar, Pascal Poupart, Jesse Hoey, Xin Jiang, and Lili Mou. 2018.
\newblock Affective neural response generation.
\newblock In \emph{European Conference on Information Retrieval}, pages
  154--166. Springer.

\bibitem[{Baheti et~al.(2018)Baheti, Ritter, Li, and
  Dolan}]{baheti2018generating}
Ashutosh Baheti, Alan Ritter, Jiwei Li, and Bill Dolan. 2018.
\newblock Generating more interesting responses in neural conversation models
  with distributional constraints.
\newblock \emph{arXiv preprint arXiv:1809.01215}.

\bibitem[{Bosselut et~al.(2018)Bosselut, Celikyilmaz, He, Gao, Huang, and
  Choi}]{bosselut-etal-2018-discourse}
Antoine Bosselut, Asli Celikyilmaz, Xiaodong He, Jianfeng Gao, Po-Sen Huang,
  and Yejin Choi. 2018.
\newblock Discourse-aware neural rewards for coherent text generation.
\newblock In \emph{Proceedings of the 2018 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long Papers)}, pages 173--184, New Orleans,
  Louisiana. Association for Computational Linguistics.

\bibitem[{Bradski(2000)}]{opencv_library}
G.~Bradski. 2000.
\newblock {The OpenCV Library}.
\newblock \emph{Dr. Dobb's Journal of Software Tools}.

\bibitem[{COLBY(1975)}]{COLBY197537}
KENNETH~MARK COLBY. 1975.
\newblock Chapter 4 - language-recognition processes for understanding
  dialogues in teletyped psychiatric interviews.
\newblock In KENNETH~MARK COLBY, editor, \emph{Artificial Paranoia}, pages 37
  -- 49. Pergamon.

\bibitem[{Danescu-Niculescu-Mizil and
  Lee(2011)}]{Danescu-Niculescu-Mizil+Lee:11a}
Cristian Danescu-Niculescu-Mizil and Lillian Lee. 2011.
\newblock Chameleons in imagined conversations: A new approach to understanding
  coordination of linguistic style in dialogs.
\newblock In \emph{Proceedings of the Workshop on Cognitive Modeling and
  Computational Linguistics, ACL 2011}.

\bibitem[{Das et~al.(2017{\natexlab{a}})Das, Kottur, Gupta, Singh, Yadav,
  Moura, Parikh, and Batra}]{das2017visual}
Abhishek Das, Satwik Kottur, Khushi Gupta, Avi Singh, Deshraj Yadav, José
  M.~F. Moura, Devi Parikh, and Dhruv Batra. 2017{\natexlab{a}}.
\newblock Visual dialog.

\bibitem[{Das et~al.(2017{\natexlab{b}})Das, Kottur, Moura, Lee, and
  Batra}]{das2017learning}
Abhishek Das, Satwik Kottur, Jos{\'e}~MF Moura, Stefan Lee, and Dhruv Batra.
  2017{\natexlab{b}}.
\newblock Learning cooperative visual dialog agents with deep reinforcement
  learning.
\newblock In \emph{Proceedings of the IEEE international conference on computer
  vision}, pages 2951--2960.

\bibitem[{Dodge et~al.(2016)Dodge, Gane, Zhang, Bordes, Chopra, Miller, Szlam,
  and Weston}]{dodge2016evaluating}
Jesse Dodge, Andreea Gane, Xiang Zhang, Antoine Bordes, Sumit Chopra, Alexander
  Miller, Arthur Szlam, and Jason Weston. 2016.
\newblock Evaluating prerequisite qualities for learning end-to-end dialog
  systems.

\bibitem[{Gao et~al.(2018)Gao, Galley, and Li}]{gao2018neural}
Jianfeng Gao, Michel Galley, and Lihong Li. 2018.
\newblock Neural approaches to conversational ai.
\newblock In \emph{The 41st International ACM SIGIR Conference on Research \&
  Development in Information Retrieval}, pages 1371--1374.

\bibitem[{Ghandeharioun et~al.(2019)Ghandeharioun, Shen, Jaques, Ferguson,
  Jones, Lapedriza, and Picard}]{ghandeharioun2019approximating}
Asma Ghandeharioun, Judy~Hanwen Shen, Natasha Jaques, Craig Ferguson, Noah
  Jones, Agata Lapedriza, and Rosalind Picard. 2019.
\newblock Approximating interactive human evaluation with self-play for
  open-domain dialog systems.
\newblock In \emph{Advances in Neural Information Processing Systems}, pages
  13658--13669.

\bibitem[{Ghazvininejad et~al.(2017)Ghazvininejad, Brockett, Chang, Dolan, Gao,
  Yih, and Galley}]{ghazvininejad2017knowledge}
Marjan Ghazvininejad, Chris Brockett, Ming-Wei Chang, Bill Dolan, Jianfeng Gao,
  Wen-tau Yih, and Michel Galley. 2017.
\newblock A knowledge-grounded neural conversation model.
\newblock \emph{arXiv preprint arXiv:1702.01932}.

\bibitem[{Guo et~al.(2019)Guo, Wang, and Wang}]{guo2019dual}
Dan Guo, Hui Wang, and Meng Wang. 2019.
\newblock Dual visual attention network for visual dialog.
\newblock In \emph{IJCAI}, pages 4989--4995.

\bibitem[{Han et~al.(2020{\natexlab{a}})Han, Meng, Wu, and Li}]{han2020non}
Qinghong Han, Yuxian Meng, Fei Wu, and Jiwei Li. 2020{\natexlab{a}}.
\newblock Non-autoregressive neural dialogue generation.
\newblock \emph{arXiv preprint arXiv:2002.04250}.

\bibitem[{Han et~al.(2020{\natexlab{b}})Han, Wallace, and
  Tsvetkov}]{han2020explaining}
Xiaochuang Han, Byron~C Wallace, and Yulia Tsvetkov. 2020{\natexlab{b}}.
\newblock Explaining black box predictions and unveiling data artifacts through
  influence functions.
\newblock \emph{arXiv preprint arXiv:2005.06676}.

\bibitem[{He et~al.(2016)He, Zhang, Ren, and Sun}]{kaiming2016resnet}
Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun. 2016.
\newblock Deep residual learning for image recognition.
\newblock In \emph{Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 770--778.

\bibitem[{He et~al.(2020)He, Yang, Yan, Li, Shen, and
  Xu}]{he-etal-2020-amalgamating}
Wanwei He, Min Yang, Rui Yan, Chengming Li, Ying Shen, and Ruifeng Xu. 2020.
\newblock Amalgamating knowledge from two teachers for task-oriented dialogue
  system with adversarial training.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 3498--3507, Online. Association
  for Computational Linguistics.

\bibitem[{Hori et~al.(2018)Hori, Alamri, Wang, Wichern, Hori, Cherian, Marks,
  Cartillier, Lopes, Das, Essa, Batra, and Parikh}]{hori2018endtoend}
Chiori Hori, Huda Alamri, Jue Wang, Gordon Wichern, Takaaki Hori, Anoop
  Cherian, Tim~K. Marks, Vincent Cartillier, Raphael~Gontijo Lopes, Abhishek
  Das, Irfan Essa, Dhruv Batra, and Devi Parikh. 2018.
\newblock End-to-end audio visual scene-aware dialog using multimodal
  attention-based video features.

\bibitem[{Hua et~al.(2020)Hua, Feng, Tao, Yan, and
  Zhang}]{10.1145/3340531.3411967}
Kai Hua, Zhiyuan Feng, Chongyang Tao, Rui Yan, and Lu~Zhang. 2020.
\newblock Learning to detect relevant contexts and knowledge for response
  selection in retrieval-based dialogue systems.
\newblock In \emph{Proceedings of the 29th ACM International Conference on
  Information Knowledge Management}, CIKM '20, page 525–534, New York, NY,
  USA. Association for Computing Machinery.

\bibitem[{Huang et~al.(2020)Huang, Zhu, and Gao}]{huang2020challenges}
Minlie Huang, Xiaoyan Zhu, and Jianfeng Gao. 2020.
\newblock Challenges in building intelligent open-domain dialog systems.
\newblock \emph{ACM Transactions on Information Systems (TOIS)}, 38(3):1--32.

\bibitem[{Jiang et~al.(2020{\natexlab{a}})Jiang, Du, Qin, Sun, and
  Yu}]{jiang2020kbgn}
Xiaoze Jiang, Siyi Du, Zengchang Qin, Yajing Sun, and Jing Yu.
  2020{\natexlab{a}}.
\newblock Kbgn: Knowledge-bridge graph network for adaptive vision-text
  reasoning in visual dialogue.
\newblock In \emph{Proceedings of the 28th ACM International Conference on
  Multimedia}, pages 1265--1273.

\bibitem[{Jiang et~al.(2020{\natexlab{b}})Jiang, Yu, Sun, Qin, Zhu, Hu, and
  Wu}]{jiang2020dam}
Xiaoze Jiang, Jing Yu, Yajing Sun, Zengchang Qin, Zihao Zhu, Yue Hu, and Qi~Wu.
  2020{\natexlab{b}}.
\newblock Dam: Deliberation, abandon and memory networks for generating
  detailed and non-repetitive responses in visual dialogue.

\bibitem[{Jiang and Bansal(2019)}]{jiang-bansal-2019-self}
Yichen Jiang and Mohit Bansal. 2019.
\newblock Self-assembling modular networks for interpretable multi-hop
  reasoning.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, pages 4474--4484, Hong Kong,
  China. Association for Computational Linguistics.

\bibitem[{Kang et~al.(2019)Kang, Lim, and Zhang}]{kang2019dual}
Gi-Cheon Kang, Jaeseo Lim, and Byoung-Tak Zhang. 2019.
\newblock Dual attention networks for visual reference resolution in visual
  dialog.
\newblock \emph{arXiv preprint arXiv:1902.09368}.

\bibitem[{Kottur et~al.(2018)Kottur, Moura, Parikh, Batra, and
  Rohrbach}]{kottur2018visual}
Satwik Kottur, Jos{\'e}~MF Moura, Devi Parikh, Dhruv Batra, and Marcus
  Rohrbach. 2018.
\newblock Visual coreference resolution in visual dialog using neural module
  networks.
\newblock In \emph{Proceedings of the European Conference on Computer Vision
  (ECCV)}, pages 153--169.

\bibitem[{Kottur et~al.(2019)Kottur, Moura, Parikh, Batra, and
  Rohrbach}]{kottur2019clevrdialog}
Satwik Kottur, José M.~F. Moura, Devi Parikh, Dhruv Batra, and Marcus
  Rohrbach. 2019.
\newblock Clevr-dialog: A diagnostic dataset for multi-round reasoning in
  visual dialog.

\bibitem[{Krishna et~al.(2017)Krishna, Zhu, Groth, Johnson, Hata, Kravitz,
  Chen, Kalantidis, Li, Shamma et~al.}]{krishna2017visual}
Ranjay Krishna, Yuke Zhu, Oliver Groth, Justin Johnson, Kenji Hata, Joshua
  Kravitz, Stephanie Chen, Yannis Kalantidis, Li-Jia Li, David~A Shamma, et~al.
  2017.
\newblock Visual genome: Connecting language and vision using crowdsourced
  dense image annotations.
\newblock \emph{International journal of computer vision}, 123(1):32--73.

\bibitem[{Krizhevsky et~al.(2012)Krizhevsky, Sutskever, and
  Hinton}]{krizhevsky2012imagenet}
Alex Krizhevsky, Ilya Sutskever, and Geoffrey~E Hinton. 2012.
\newblock Imagenet classification with deep convolutional neural networks.
\newblock In \emph{Advances in neural information processing systems}, pages
  1097--1105.

\bibitem[{Lee et~al.(2019)Lee, Zhu, Takanobu, Li, Zhang, Zhang, Li, Peng, Li,
  Huang et~al.}]{lee2019convlab}
Sungjin Lee, Qi~Zhu, Ryuichi Takanobu, Xiang Li, Yaoqin Zhang, Zheng Zhang,
  Jinchao Li, Baolin Peng, Xiujun Li, Minlie Huang, et~al. 2019.
\newblock Convlab: Multi-domain end-to-end dialog system platform.
\newblock \emph{arXiv preprint arXiv:1904.08637}.

\bibitem[{Lewis et~al.(2017)Lewis, Yarats, Dauphin, Parikh, and
  Batra}]{lewis-etal-2017-deal}
Mike Lewis, Denis Yarats, Yann Dauphin, Devi Parikh, and Dhruv Batra. 2017.
\newblock Deal or no deal? end-to-end learning of negotiation dialogues.
\newblock In \emph{Proceedings of the 2017 Conference on Empirical Methods in
  Natural Language Processing}, pages 2443--2453, Copenhagen, Denmark.
  Association for Computational Linguistics.

\bibitem[{Li(2020)}]{li2020teaching}
Jiwei Li. 2020.
\newblock Teaching machines to converse.
\newblock \emph{arXiv preprint arXiv:2001.11701}.

\bibitem[{Li et~al.(2015)Li, Galley, Brockett, Gao, and
  Dolan}]{li2015diversity}
Jiwei Li, Michel Galley, Chris Brockett, Jianfeng Gao, and Bill Dolan. 2015.
\newblock A diversity-promoting objective function for neural conversation
  models.
\newblock \emph{arXiv preprint arXiv:1510.03055}.

\bibitem[{Li et~al.(2016{\natexlab{a}})Li, Galley, Brockett, Spithourakis, Gao,
  and Dolan}]{li2016persona}
Jiwei Li, Michel Galley, Chris Brockett, Georgios~P Spithourakis, Jianfeng Gao,
  and Bill Dolan. 2016{\natexlab{a}}.
\newblock A persona-based neural conversation model.
\newblock \emph{arXiv preprint arXiv:1603.06155}.

\bibitem[{Li et~al.(2016{\natexlab{b}})Li, Monroe, Ritter, Galley, Gao, and
  Jurafsky}]{li2016deep}
Jiwei Li, Will Monroe, Alan Ritter, Michel Galley, Jianfeng Gao, and Dan
  Jurafsky. 2016{\natexlab{b}}.
\newblock Deep reinforcement learning for dialogue generation.
\newblock \emph{arXiv preprint arXiv:1606.01541}.

\bibitem[{Li et~al.(2017)Li, Monroe, Shi, Jean, Ritter, and
  Jurafsky}]{li2017adversarial}
Jiwei Li, Will Monroe, Tianlin Shi, S{\'e}bastien Jean, Alan Ritter, and Dan
  Jurafsky. 2017.
\newblock Adversarial learning for neural dialogue generation.
\newblock \emph{arXiv preprint arXiv:1701.06547}.

\bibitem[{Lison and Tiedemann(2016)}]{Lison2016OpenSubtitles2016EL}
P.~Lison and J.~Tiedemann. 2016.
\newblock Opensubtitles2016: Extracting large parallel corpora from movie and
  tv subtitles.
\newblock In \emph{LREC}.

\bibitem[{Lowe et~al.(2015)Lowe, Pow, Serban, and Pineau}]{lowe2015ubuntu}
Ryan Lowe, Nissan Pow, Iulian Serban, and Joelle Pineau. 2015.
\newblock The ubuntu dialogue corpus: A large dataset for research in
  unstructured multi-turn dialogue systems.
\newblock \emph{arXiv preprint arXiv:1506.08909}.

\bibitem[{Lu et~al.(2017)Lu, Kannan, Yang, Parikh, and Batra}]{lu2017best}
Jiasen Lu, Anitha Kannan, Jianwei Yang, Devi Parikh, and Dhruv Batra. 2017.
\newblock Best of both worlds: Transferring knowledge from discriminative
  learning to a generative visual dialog model.

\bibitem[{Massiceti et~al.(2018)Massiceti, Siddharth, Dokania, and
  Torr}]{massiceti2018flipdial}
Daniela Massiceti, N~Siddharth, Puneet~K Dokania, and Philip~HS Torr. 2018.
\newblock Flipdial: A generative model for two-way visual dialogue.
\newblock In \emph{Proceedings of the IEEE Conference on Computer Vision and
  Pattern Recognition}, pages 6097--6105.

\bibitem[{Meng et~al.(2020)Meng, Wang, Han, Sun, Wu, Yan, and
  Li}]{meng2020openvidial}
Yuxian Meng, Shuhe Wang, Qinghong Han, Xiaofei Sun, Fei Wu, Rui Yan, and Jiwei
  Li. 2020.
\newblock Openvidial: A large-scale, open-domain dialogue dataset with visual
  contexts.
\newblock \emph{arXiv preprint arXiv:2012.15015}.

\bibitem[{Niu et~al.(2019)Niu, Zhang, Zhang, Zhang, Lu, and
  Wen}]{niu2019recursive}
Yulei Niu, Hanwang Zhang, Manli Zhang, Jianhong Zhang, Zhiwu Lu, and Ji-Rong
  Wen. 2019.
\newblock Recursive visual attention in visual dialog.
\newblock In \emph{Proceedings of the IEEE Conference on Computer Vision and
  Pattern Recognition}, pages 6679--6688.

\bibitem[{Papineni et~al.(2002)Papineni, Roukos, Ward, and
  Zhu}]{papineni2002bleu}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002.
\newblock Bleu: a method for automatic evaluation of machine translation.
\newblock In \emph{Proceedings of the 40th annual meeting of the Association
  for Computational Linguistics}, pages 311--318.

\bibitem[{Park et~al.(2020)Park, Whang, Yoon, and Lim}]{park2020multi}
Sungjin Park, Taesun Whang, Yeochan Yoon, and Hueiseok Lim. 2020.
\newblock Multi-view attention networks for visual dialog.
\newblock \emph{arXiv preprint arXiv:2004.14025}.

\bibitem[{Rashkin et~al.(2018)Rashkin, Smith, Li, and
  Boureau}]{rashkin2018towards}
Hannah Rashkin, Eric~Michael Smith, Margaret Li, and Y-Lan Boureau. 2018.
\newblock Towards empathetic open-domain conversation models: A new benchmark
  and dataset.
\newblock \emph{arXiv preprint arXiv:1811.00207}.

\bibitem[{Ren et~al.(2015)Ren, He, Girshick, and Sun}]{ren2015faster}
Shaoqing Ren, Kaiming He, Ross Girshick, and Jian Sun. 2015.
\newblock Faster r-cnn: Towards real-time object detection with region proposal
  networks.
\newblock In \emph{Advances in neural information processing systems}, pages
  91--99.

\bibitem[{Roller et~al.(2020)Roller, Dinan, Goyal, Ju, Williamson, Liu, Xu,
  Ott, Shuster, Smith et~al.}]{roller2020recipes}
Stephen Roller, Emily Dinan, Naman Goyal, Da~Ju, Mary Williamson, Yinhan Liu,
  Jing Xu, Myle Ott, Kurt Shuster, Eric~M Smith, et~al. 2020.
\newblock Recipes for building an open-domain chatbot.
\newblock \emph{arXiv preprint arXiv:2004.13637}.

\bibitem[{Sennrich et~al.(2016)Sennrich, Haddow, and
  Birch}]{sennrich-etal-2016-neural}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016.
\newblock Neural machine translation of rare words with subword units.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 1715--1725,
  Berlin, Germany. Association for Computational Linguistics.

\bibitem[{Seo et~al.(2017)Seo, Lehrmann, Han, and Sigal}]{seo2017visual}
Paul~Hongsuck Seo, Andreas Lehrmann, Bohyung Han, and Leonid Sigal. 2017.
\newblock Visual reference resolution using attention memory for visual dialog.
\newblock In \emph{Advances in neural information processing systems}, pages
  3719--3729.

\bibitem[{Serban et~al.(2016)Serban, Sordoni, Lowe, Charlin, Pineau, Courville,
  and Bengio}]{serban2016hierarchical}
Iulian~Vlad Serban, Alessandro Sordoni, Ryan Lowe, Laurent Charlin, Joelle
  Pineau, Aaron Courville, and Yoshua Bengio. 2016.
\newblock A hierarchical latent variable encoder-decoder model for generating
  dialogues.
\newblock \emph{arXiv preprint arXiv:1605.06069}.

\bibitem[{Shao et~al.(2017)Shao, Gouws, Britz, Goldie, Strope, and
  Kurzweil}]{shao2017generating}
Louis Shao, Stephan Gouws, Denny Britz, Anna Goldie, Brian Strope, and Ray
  Kurzweil. 2017.
\newblock Generating high-quality and informative conversation responses with
  sequence-to-sequence models.
\newblock \emph{arXiv preprint arXiv:1701.03185}.

\bibitem[{Sordoni et~al.(2015)Sordoni, Galley, Auli, Brockett, Ji, Mitchell,
  Nie, Gao, and Dolan}]{sordoni2015neural}
Alessandro Sordoni, Michel Galley, Michael Auli, Chris Brockett, Yangfeng Ji,
  Margaret Mitchell, Jian-Yun Nie, Jianfeng Gao, and Bill Dolan. 2015.
\newblock A neural network approach to context-sensitive generation of
  conversational responses.
\newblock \emph{arXiv preprint arXiv:1506.06714}.

\bibitem[{Tian et~al.(2017)Tian, Yan, Mou, Song, Feng, and
  Zhao}]{tian-etal-2017-make}
Zhiliang Tian, Rui Yan, Lili Mou, Yiping Song, Yansong Feng, and Dongyan Zhao.
  2017.
\newblock How to make context more useful? an empirical study on context-aware
  neural conversational models.
\newblock In \emph{Proceedings of the 55th Annual Meeting of the Association
  for Computational Linguistics (Volume 2: Short Papers)}, pages 231--236,
  Vancouver, Canada. Association for Computational Linguistics.

\bibitem[{Tiedemann(2009)}]{Tiedemann2009NewsFO}
J.~Tiedemann. 2009.
\newblock News from opus — a collection of multilingual parallel corpora with
  tools and interfaces.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{vaswani2017attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, {\L}ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock Attention is all you need.
\newblock In \emph{Advances in neural information processing systems}, pages
  5998--6008.

\bibitem[{Vinyals and Le(2015)}]{vinyals2015neural}
Oriol Vinyals and Quoc Le. 2015.
\newblock A neural conversational model.
\newblock \emph{arXiv preprint arXiv:1506.05869}.

\bibitem[{de~Vries et~al.(2017)de~Vries, Strub, Chandar, Pietquin, Larochelle,
  and Courville}]{devries2017guesswhat}
Harm de~Vries, Florian Strub, Sarath Chandar, Olivier Pietquin, Hugo
  Larochelle, and Aaron Courville. 2017.
\newblock Guesswhat?! visual object discovery through multi-modal dialogue.

\bibitem[{Wallace(2009)}]{Wallace2009}
Richard~S. Wallace. 2009.
\newblock \emph{The Anatomy of A.L.I.C.E.}, pages 181--210. Springer
  Netherlands, Dordrecht.

\bibitem[{Wang et~al.(2021)Wang, Meng, Sun, Wu, Ouyang, Yan, Zhang, and
  Li}]{wang2021modeling}
Shuhe Wang, Yuxian Meng, Xiaofei Sun, Fei Wu, Rongbin Ouyang, Rui Yan, Tianwei
  Zhang, and Jiwei Li. 2021.
\newblock Modeling text-visual mutual dependency for multi-modal dialog
  generation.
\newblock \emph{arXiv preprint arXiv:2105.14445}.

\bibitem[{Weizenbaum(1966)}]{weizenbaum1966eliza}
Joseph Weizenbaum. 1966.
\newblock Eliza—a computer program for the study of natural language
  communication between man and machine.
\newblock \emph{Communications of the ACM}, 9(1):36--45.

\bibitem[{Wu et~al.(2018)Wu, Wang, Shen, Reid, and Van Den~Hengel}]{wu2018you}
Qi~Wu, Peng Wang, Chunhua Shen, Ian Reid, and Anton Van Den~Hengel. 2018.
\newblock Are you talking to me? reasoned visual dialog generation through
  adversarial learning.
\newblock In \emph{Proceedings of the IEEE Conference on Computer Vision and
  Pattern Recognition}, pages 6106--6115.

\bibitem[{Xie et~al.(2017)Xie, Wang, Li, L{\'e}vy, Nie, Jurafsky, and
  Ng}]{xie2017data}
Ziang Xie, Sida~I Wang, Jiwei Li, Daniel L{\'e}vy, Aiming Nie, Dan Jurafsky,
  and Andrew~Y Ng. 2017.
\newblock Data noising as smoothing in neural network language models.
\newblock \emph{arXiv preprint arXiv:1703.02573}.

\bibitem[{Xu et~al.(2018)Xu, Ren, Lin, and Sun}]{xu2018dp}
Jingjing Xu, Xuancheng Ren, Junyang Lin, and Xu~Sun. 2018.
\newblock Dp-gan: diversity-promoting generative adversarial network for
  generating informative and diversified text.
\newblock \emph{arXiv preprint arXiv:1802.01345}.

\bibitem[{Yang et~al.(2019)Yang, Zha, and Zhang}]{yang2019making}
Tianhao Yang, Zheng-Jun Zha, and Hanwang Zhang. 2019.
\newblock Making history matter: History-advantage sequence training for visual
  dialog.

\bibitem[{Young et~al.(2017)Young, Cambria, Chaturvedi, Huang, Zhou, and
  Biswas}]{young2017augmenting}
Tom Young, Erik Cambria, Iti Chaturvedi, Minlie Huang, Hao Zhou, and Subham
  Biswas. 2017.
\newblock Augmenting end-to-end dialog systems with commonsense knowledge.
\newblock \emph{arXiv preprint arXiv:1709.05453}.

\bibitem[{Zhang et~al.(2018)Zhang, Dinan, Urbanek, Szlam, Kiela, and
  Weston}]{zhang2018personalizing}
Saizheng Zhang, Emily Dinan, Jack Urbanek, Arthur Szlam, Douwe Kiela, and Jason
  Weston. 2018.
\newblock Personalizing dialogue agents: I have a dog, do you have pets too?
\newblock \emph{arXiv preprint arXiv:1801.07243}.

\bibitem[{Zhang et~al.(2019)Zhang, Sun, Galley, Chen, Brockett, Gao, Gao, Liu,
  and Dolan}]{zhang2019dialogpt}
Yizhe Zhang, Siqi Sun, Michel Galley, Yen-Chun Chen, Chris Brockett, Xiang Gao,
  Jianfeng Gao, Jingjing Liu, and Bill Dolan. 2019.
\newblock Dialogpt: Large-scale generative pre-training for conversational
  response generation.
\newblock \emph{arXiv preprint arXiv:1911.00536}.

\bibitem[{Zhao et~al.(2019)Zhao, Xie, and Eskenazi}]{zhao2019rethinking}
Tiancheng Zhao, Kaige Xie, and Maxine Eskenazi. 2019.
\newblock Rethinking action spaces for reinforcement learning in end-to-end
  dialog agents with latent variable models.
\newblock \emph{arXiv preprint arXiv:1902.08858}.

\bibitem[{Zhao et~al.(2017)Zhao, Zhao, and Eskenazi}]{zhao2017learning}
Tiancheng Zhao, Ran Zhao, and Maxine Eskenazi. 2017.
\newblock Learning discourse-level diversity for neural dialog models using
  conditional variational autoencoders.
\newblock \emph{arXiv preprint arXiv:1703.10960}.

\bibitem[{Zhao et~al.(2020)Zhao, Wu, Xu, Tao, Zhao, and
  Yan}]{zhao-etal-2020-knowledge-grounded}
Xueliang Zhao, Wei Wu, Can Xu, Chongyang Tao, Dongyan Zhao, and Rui Yan. 2020.
\newblock Knowledge-grounded dialogue generation with pre-trained language
  models.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 3377--3390, Online. Association
  for Computational Linguistics.

\bibitem[{Zheng et~al.(2019)Zheng, Wang, Qi, and Zhu}]{zheng2019reasoning}
Zilong Zheng, Wenguan Wang, Siyuan Qi, and Song-Chun Zhu. 2019.
\newblock Reasoning visual dialogs with structural and partial observations.
\newblock In \emph{Proceedings of the IEEE Conference on Computer Vision and
  Pattern Recognition}, pages 6669--6678.

\bibitem[{Zhou et~al.(2017)Zhou, Huang, Zhang, Zhu, and
  Liu}]{zhou2017emotional}
Hao Zhou, Minlie Huang, Tianyang Zhang, Xiaoyan Zhu, and Bing Liu. 2017.
\newblock Emotional chatting machine: Emotional conversation generation with
  internal and external memory.
\newblock \emph{arXiv preprint arXiv:1704.01074}.

\bibitem[{Zhou et~al.(2020)Zhou, Gao, Li, and Shum}]{zhou2020design}
Li~Zhou, Jianfeng Gao, Di~Li, and Heung-Yeung Shum. 2020.
\newblock The design and implementation of xiaoice, an empathetic social
  chatbot.
\newblock \emph{Computational Linguistics}, 46(1):53--93.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{emnlp2020}

\usepackage[utf8]{inputenc} % allow utf-8 input
\usepackage{microtype}
\usepackage{graphicx}
\usepackage{subfigure}
\usepackage{booktabs} 
\usepackage{comment}
\usepackage{hyperref}
\usepackage{mathtools}
\usepackage{color}
\usepackage{graphicx}
\usepackage{pifont}
\usepackage{CJKutf8}
\usepackage{url}
\usepackage{multicol,multirow}
\usepackage{makecell}
\usepackage{array}
\usepackage{bm}
\usepackage{parskip}
\usepackage[T1]{fontenc}    % use 8-bit T1 fonts
\usepackage{booktabs}       % professional-quality tables
\usepackage{amsmath, amsthm}
\usepackage{amssymb}
\usepackage{amsfonts}

\usepackage{times}
\usepackage{float}
\usepackage{stfloats}
\usepackage{latexsym}

\usepackage{microtype}
\definecolor{ao}{rgb}{0.0, 0.5, 0.0}
\definecolor{asparagus}{rgb}{0.53, 0.66, 0.42}
\definecolor{amber}{rgb}{1.0, 0.49, 0.0}
\definecolor{alizarin}{rgb}{0.82, 0.1, 0.26}
\definecolor{applegreen}{rgb}{0.55, 0.71, 0.0}
\definecolor{amethyst}{rgb}{0.6, 0.4, 0.8}
\definecolor{auburn}{rgb}{0.43, 0.21, 0.1}
\aclfinalcopy % Uncomment this line for the final submission
\title{OpenViDial 2.0: A Larger-Scale, Open-Domain Dialogue Generation\\Dataset with Visual Contexts}
\date{}

\author{
Shuhe Wang$^{\clubsuit\spadesuit}$, 
Yuxian Meng$^\clubsuit$, 
Xiaoya Li$^\clubsuit$\\
{\bf Xiaofei Sun$^\clubsuit$,
Rongbin Ouyang$^\spadesuit$
Jiwei Li$^{\blacklozenge\clubsuit}$}\\
$^\clubsuit$ Shannon.AI, $^\spadesuit$Peking University, $^\blacklozenge$Zhejiang University\\ 
  \{shuhe\_wang, yuxian\_meng, xiaoya\_li, xiaofei\_sun, jiwei\_li\}@shannonai.com\\

}

\begin{document}
\maketitle

\begin{abstract}
In order to better simulate the real human conversation process, models need to generate dialogue utterances based on not only preceding textual contexts but also visual contexts. However, with the development of multi-modal dialogue learning, the dataset scale gradually becomes a bottleneck.
In this report, we release {\bf OpenViDial 2.0}, a larger-scale open-domain multi-modal dialogue dataset compared to the previous version OpenViDial 1.0 \citep{meng2020openvidial}. OpenViDial 2.0 contains a total number of 5.6 million dialogue turns extracted from either movies or TV series from different resources, and each dialogue turn is paired with its corresponding visual context.
We hope this large-scale dataset can help facilitate future researches on open-domain multi-modal dialog generation, e.g., multi-modal pretraining for dialogue generation.\footnote{Dataset is available found at 
{\url{https://github.com/ShannonAI/OpenViDial}}.}

\end{abstract}

\section{Introduction}

Developing open-domain dialogue agents is of growing interest \citep{,li2017adversarial,ghazvininejad2017knowledge,zhou2017emotional,gao2018neural,asghar2018affective,han2020non,zhou2020design}. Existing methods for developing effective open-domain dialogue agents mostly follow a two-step pipeline: (1) collecting a large-scale dataset containing massive dialog turns from real conversations, and (2) training a neural model to learn to generate high quality responses given the previous dialogue contexts \cite{li2016deep,li2016persona,zhang2018personalizing,huang2020challenges}.

Since most methods are data-driven, a large-scale and high quality open-domain dialogue datasets may be the first matter to be considered before designing the model. \citet{meng2020openvidial} released the OpenViDial dataset which contains a total number of 1.1 million dialogue turns with utterances paired with visual context. Some recent works leveraged the OpenViDial dataset and built effective multi-modal dialog models \citep{wang2021modeling} on top, demonstrating that learning multi-modal features gives rise to higher response quality.

In this report, we collect and extend OpenViDial, releasing OpenViDial 2.0, a much larger-scale open-domain dialogue dataset with visual contexts. In common with the prior version OpenViDial 1.0 \citep{meng2020openvidial}, the dialogue turns and visual contexts in OpenViDial 2.0 are also extracted from movies and TV series, where each dialogue turn is paired with the corresponding visual context in which it takes place. OpenViDial 2.0 contains a total number of 5.6 million dialogue turns along with 5.6 million visual contexts stored as images, a scale of 4 times larger than OpenViDial 1.0. 
We hope this large-scale dataset can help facilitate future researches on open-domain multi-modal dialog generation, e.g., multi-modal pretraining for dialogue generation. 

\section{Related Work}
\subsection{Open Domain Dialog Datasets}

\paragraph{Textual Dialog Datasets} Since the task of open-domain dialog generation has developed for many years, there are various open-domain dialog datasets only consists textual information. For simulating the movie conversation, there are OpenSubtitle dataset \citep{Tiedemann2009NewsFO, Lison2016OpenSubtitles2016EL} and Cornell Movie-Dialogs Corpus \citep{Danescu-Niculescu-Mizil+Lee:11a}. The OpenSubtitle dataset is a large-scale dataset contains a total number of 3.35G sentence fragments extracted from the OpenSubtitle website, while the Cornell Movie-Dialogs Corpus contains a collection of movie conversations extracted from raw movie scripts. For simulating the social conversation, there are PersonaChat \cite{zhang2018personalizing} and Twitter Triple Corpus \cite{sordoni2015neural}. The Twitter Triple Corpus consists of 4,232 Twitter conversation triples evaluated from 33K candidate triples by human raters. Other datasets such as the Ubuntu Dialog Corpus \cite{lowe2015ubuntu} and EmpatheticDialogues \cite{rashkin2018towards} are both commonly used for textual open-domain dialog generation.

\paragraph{Visual Dialog Datasets} A mount of datasets containing visual features have been developed, since the task of VisualDialog is first introduced by \citet{das2017visual}, where a model is required to answer questions by given a dialog history and the image itself as contexts. For this work, \citet{das2017visual} released VisDial v0.9 and v1.0 datasets which contains 120K images from MSCOCO\footnote{\url{http://mscoco.org/}} and each image is associated with 10 rounds of question-answer dialog. Further, other datasets like the GuessWhat?! dataset \cite{devries2017guesswhat}, the CLEVERDialog dataset \cite{kottur2019clevrdialog}, the MNIST-Dialog dataset \cite{seo2017visual} and the Audio Visual Scene-Aware Dialog (AVSD) dataset \citep{hori2018endtoend,alamri2019audio} are mainly focus more on answering questions according to an image or video rather than dialogue generation with visual contexts.The OpenViDial dataset \cite{meng2020openvidial} is released to alleviate this situation, where contains 1.1M dialogue turns and each dialogue turn paired with the corresponding visual context in which it takes place. And thus, models need to learn to generate dialogue utterances not only based on preceding textual contexts but also visual contexts.

\subsection{Dialog Generation}

\paragraph{Open Domain Dialog Generation} Open-domain dialog generation is a simulation for real human conversations and is a traditional task in NLP \citep{weizenbaum1966eliza, COLBY197537, Wallace2009}. Currently, the most researches for open-domain dialog generation are based on sequence-to-sequence architecture \citep{vinyals2015neural, li2015diversity, dodge2016evaluating, serban2016hierarchical, zhao2017learning, xie2017data, lee2019convlab, ghandeharioun2019approximating, li2020teaching, han2020explaining, zhang2019dialogpt, roller2020recipes}. And whether a model can generate diverse \citep{xu2018dp, baheti2018generating}, coherent \citep{li2016deep,li2017adversarial,tian-etal-2017-make,bosselut-etal-2018-discourse,adiwardana2020towards}, informative \citep{shao2017generating,lewis-etal-2017-deal,ghazvininejad2017knowledge,young2017augmenting,zhao2019rethinking} and knowledge-fused \citep{10.1145/3340531.3411967,zhao-etal-2020-knowledge-grounded,he-etal-2020-amalgamating} responses or not has become metrics to evaluate a dialog generation model. However, the mainly researches described above are developed on textual only and the development of multi-modal dialog generation is relatively slow since the lack of large-scale datasets.

\paragraph{Visual Dialog Generation} Most of existing works apply attention mechanisms to model the interplay between text and visual contexts \citep{lu2017best, kottur2018visual, jiang-bansal-2019-self, yang2019making, guo2019dual, niu2019recursive, kang2019dual, park2020multi, jiang2020dam}. Other techniques like reinforcement learning \citep{das2017learning, wu2018you}, variational auto-encoders \cite{massiceti2018flipdial} and graph networks \citep{zheng2019reasoning, jiang2020kbgn} have also been employed to the visual dialog task. More recently, based on the OpenViDial dataset \cite{meng2020openvidial}, \citet{wang2021modeling} proposed three attention-based models \cite{vaswani2017attention} to generate dialogue utterances given the preceding text-visual contexts and further proposed to build text-visual dependency to improve the dialogue quality, making an initial step for the task of text-visual open-domain dialogue generation rather than answering questions based on an image.

\begin{table*}
\centering
\begin{tabular}{lcc} \\\toprule
\textbf{Statistics} & \textbf{OpenViDial 1.0} & \textbf{OpenViDial 2.0} \\\midrule
Number of turns & 1.1M & 5.6M \\
Number of images & 1.1M & 5.6M \\
Vocab size before BPE & 70K & 278K \\
Vocab size after BPE & 30K & 30K \\
Average length of each episode & 14 & 48 \\
Average length of each turn & 7.6 & 8.3 \\\bottomrule
\end{tabular}
\caption{Detailed statistics for OpenViDial 2.0 and a comparison to OpenViDial 1.0.}
\label{tab:stats}
\end{table*}

\begin{table}[t]
\centering
\begin{tabular}{lcc} \\\toprule
& \textbf{OpenViDial 1.0} & \textbf{OpenViDial 2.0} \\\midrule
Train & 1M & 4.6M \\
Dev & 50K & 0.5M \\
Test & 50K & 0.5M \\\bottomrule
\end{tabular}
\caption{Splitting for training, dev and test}
\label{tab:split}
\end{table}

\section{Constructing OpenViDial 2.0}
In this section, we describe the details of constructing of OpenViDial 2.0. We first collect a raw dataset consisting of about 800 English movies and TV series with an average length of 2.5 hours per video. Each video has a corresponding external English subtitle file where each line is a string including the subtitle text and the time interval. There is no video embedded with any internal subtitles.

The full process to construct OpenViDial 2.0 can be divided into three steps: (1) segmenting each video into multiple frames; (2) pairing each frame with subtitle text from it corresponding subtitle file; (3) splitting these (image, text) pairs into different dialog turns. The OpenCV \cite{opencv_library} toolkit is used to segment each video into multiple images by frame, and we discard the initial and the last 10 minutes of each video because of the general existence of intro in movies and TV series. To pair images with textual subtitles for each video, we first read the video's subtitle file row-by-row and obtain the time interval as well as the subtitle text. Then, we extract a group of images according to the time interval, and randomly choose one image from the group as the visual context paired with the subtitle text, forming a paired (image, text) dialog turn.

We are able to construct a final dataset of 5.6M dialog turns, where each turn consists of a sequence of words and an image. The size of the image is one of (1) 1280$\times$720, (2) 1920$\times$1080, and (3) 2048$\times$1080 according to different video resources. We employ the BPE tokenizer \cite{sennrich-etal-2016-neural} to preprocess the text. A detailed comparison with OpenViDial 1.0 is shown in Table \ref{tab:stats}. The splitting for training, dev and test is shown in Table \ref{tab:split}.

\begin{table*}
\centering
\small
\scalebox{0.85}{
\begin{tabular}{lcccc} \\\toprule
{\bf Dataset} & {\bf Genre} & {\bf Multi-Modal?} & {\bf \# Sentences} &  {\bf \# Images}\\\midrule 
OpenSubtitles 2016 \citep{Lison2016OpenSubtitles2016EL} & Plain-text Dialog & \ding{55} &337M & -- \\
Cornell Movie-Dialogs \citep{Danescu-Niculescu-Mizil+Lee:11a} & Plain-text Dialog & \ding{55}  &0.3M &--  \\
VisDial v1.0 \citep{das2017visual}  & VQA & \checkmark & 2.4M & 120K  \\
Guess-What?! \citep{devries2017guesswhat}  & VQA & \checkmark & 0.8M & 66K \\
AVSD \citep{alamri2019audio} &  VQA & \checkmark & 152K & -- \\
OpenViDial 1.0 \citep{meng2020openvidial} & Visual+Text Dialog &  \checkmark & 1.1M & 1.1M  \\
OpenViDial 2.0 & Visual+Text Dialog &  \checkmark & 5.6M & 5.6M  \\\bottomrule
\end{tabular}
}
\caption{A comparison of different datasets. VQA: Visual Question Answering.}
\label{tab:comparison}
\end{table*}

In Table \ref{tab:comparison}, we make a comparison with existing widely-used dialog datasets. Both OpenViDial 1.0 and OpenViDial 2.0 focus on multi-modal dialog generation in comparison to VisDial, Guess-What?! and AVSD which focus more on VQA.
Comparing against OpenViDial 1.0, OpenViDial 2.0 is much larger in scale,  about 5 times as big as OpenViDial 1.0.

\begin{table*}
  \small
  \centering
  \begin{tabular}{l|cccccc} \toprule
  \multicolumn{1}{l}{\bf System} & \multicolumn{1}{l}{\bf Model} & \multicolumn{1}{l}{\bf BLEU} & 
  \multicolumn{1}{l}{\bf Dis-1} & \multicolumn{1}{l}{\bf Dis-2} & \multicolumn{1}{l}{\bf Dis-3} & \multicolumn{1}{l}{\bf Dis-4}\\\midrule 
  \multirow{2}{*}{\bf NV} & {\it w/o MI} & \text{1.95} & \text{0.0037} & \text{0.0302} & \text{0.0929}  & \text{0.1711}\\
   & {\it w/ MI} & \text{1.96} & \text{0.0039} & \text{0.0311} & \text{0.0953}  & \text{0.1630}\\\hline
   \multirow{2}{*}{CV} & {\it w/o MI} & \text{1.97} & \text{0.0041} & \text{0.0353} & \text{0.0999} & \text{0.1726}\\
    & {\it w/ MI} & \text{1.98} & \text{0.0047} & \text{0.0392} & \text{0.1093} & \text{0.1774} \\\hline
   \multirow{2}{*}{FV} & {\it w/o MI} & \text{1.99} & \text{0.0056} & \text{0.0431} & \text{0.1250} & \text{0.2215}\\
    & {\it w/ MI} & \text{2.00} & \text{0.0060} & \text{0.0460} & \text{0.1321} & \text{0.2311}\\
  \bottomrule
  \end{tabular}
  \caption{Automatic evaluation results for BLEU, Stopword\% and Diversity. }
  \label{tab:bleu}
\end{table*}

To evaluate OpenViDial 2.0, we experiment on OpenViDial2.0 using multi-modal dialog models proposed in \cite{wang2021modeling}. 

\subsection{Vanilla Visual Dialog Models}
According to the granularity of the visual features ranges from none, coarse-grained image features to fine-grained object features, \citet{wang2021modeling} proposed three vanilla visual dialog models: (1) the NoVisual(NV) model, (2) the CoarseVisual(CV) model and (3) the FineVisual(FV) model.

\paragraph{NoVisual} The NV model is a general uni-modal dialog generation model, which is required to learn to generate responses using only dialog texts without visual information. A standard Transformer \cite{vaswani2017attention} architecture is used as the backbone for the NV model. For each dialog turn, all the preceding dialog texts are packed into a long sequence with a special token as the delimiter. Then, this sequence is embedded with positional encodings including sentence-level positional encoding and token-level positional encoding. Last, it is fed to the Transfromer as input.

\paragraph{CoarseVisual} In contrast to the NV model, the CV model injects coarse-level visual information into dialog generation. For each dialog turn, it utilizes a ResNet-50 model \cite{kaiming2016resnet} pre-trained on ImageNet \cite{krizhevsky2012imagenet} to extract a high-dimensional feature for each image as the visual information. Then the image feature is added to its corresponding text representation forming the text-visual feature. Positional encodings are also used to notify position information. The concatenated long text-visual sequence is fed into the Transformer model.

\paragraph{FineVisual} Extracting visual information from a coarse view might be insufficient to model fine-grained visual elements in images such as facial expressions, body gestures as well as physical motions. The FV model thus uses Faster R-CNN \cite{ren2015faster} pre-trained on Visual Genome \cite{krishna2017visual} to extract fine-grained visual features. Different from the CV model, the FV model directly concatenates the set of extracted fine-grained visual information with the dialog texts into a long sequence. And except for  the sentence-level and token-level positional embeddings, there is an additional positional embedding for visual features. 

\subsection{Visual-Text Mutual Dependeny}
Although each response is generated according to the preceding textual and visual contexts, there is no guarantee on whether or how much the visual contexts are used. To significantly strength the connection between the generated response and its visual contexts, \citet{wang2021modeling} proposed to model the mutual information (MI) between visual contexts and text features. To put it simply, we use visual feature to represent both the coarse-grained feature and the fine-grained feature. For building the connection between visual contexts and textual utterances, a light discriminative network is trained. The whole requirement for the discriminative network is to discriminate the degree of the connection between the given visual feature and textual feature. In each inference step, both the CV and FV model are required to generate N-best responses list with its probability as the forward probability rather than only the best response. And each response in N-best list along with the preceding visual feature are fed into the former trained discriminative network obtaining the backward probability. Finally, the forward probability and the backward probability is concatenated to rerank the N-best list. For more details please refer to \citet{wang2021modeling}.

\subsection{Results}
Following \citet{wang2021modeling}, we report the results in terms of the following automatic evaluation metrics: 
\begin{itemize}
    \item {\bf BLEU}: BLEU score is a common automatic evaluation method for majority NLP tasks \citep{papineni2002bleu, sordoni2015neural}, which score the n-gram overlaps between the generated sequences and reference sequences. For our experiment we report the BLEU-4 score.
    \item {\bf Diversity}: Diversity is usually reported in the task of dialogue generation \cite{li2015diversity}, which score the number of distinct n-grams in generated responses, and $n=1,2,3,4$ for this experiment. 
\end{itemize}
Results are shown in Table \ref{tab:bleu}. Since OpenViDial 2.0 is much larger than OpenViDial 1.0, we only use the top 5 objects for FineVisual model compared to using top 20 objects on OpenViDial 1.0, and this is the main reason why FV doesn't significantly perform better than FV and NV. 

\section{Conclusion}
In this report, we release OpenViDial 2.0, a larger-scale open-domain multi-modal dialogue dataset with visual contexts, updated from the previous version 1.0. 
OpenViDial 2.0 contains a total number of 5.6 million dialogue turns extracted from either movies or TV series from different resources,
and is four times larger than version 1.0 at scale. 
We hope this large-scale dataset can help facilitate future researches on open-domain multi-modal dialog generation.
OpenViDial 2.0 is available at \url{https://github.com/ShannonAI/OpenViDial}.

\bibliography{emnlp2020}
\bibliographystyle{acl_natbib}

\end{document}